Solve the system of linear equations using matrix inverse method.

x-y+z=1

2y-z=1

2x+3y=1


1
Expert's answer
2022-06-16T10:31:09-0400
"A=\begin{pmatrix}
 1 & -1 & 1 \\
 0 & 2 & -1 \\ 
 2 & 3 & 0 \\
\end{pmatrix}, X=\begin{pmatrix}
 x \\ 
 y\\
z
\end{pmatrix}, B=\begin{pmatrix}
 1\\ 
1\\
1
\end{pmatrix}"

"AX=B"

"A^{-1}AX=A^{-1}B=>X=A^{-1}B"

augment the matrix with the identity matrix:


"\begin{pmatrix}
 1 & -1 & 1 & & 1& 0 & 0 \\
 0 & 2 & -1 & & 0& 1 & 0 \\ 
 2 & 3 & 0 & & 0 & 0 & 1 \\
\end{pmatrix}"

"R_3=R_3-2R_1"


"\begin{pmatrix}
 1 & -1 & 1 & & 1& 0 & 0 \\
 0 & 2 & -1 & & 0& 1 & 0 \\ 
 0 & 5 & -2 & & -2 & 0 & 1 \\
\end{pmatrix}"

"R_2=R_2\/2"


"\begin{pmatrix}
 1 & -1 & 1 & & 1& 0 & 0 \\
 0 & 1 & -1/2 & & 0& 1/2 & 0 \\ 
 0 & 5 & -2 & & -2 & 0 & 1 \\
\end{pmatrix}"

"R_1=R_1+R_2"


"\begin{pmatrix}
 1 & 0 & 1/2 & & 1& 1/2 & 0 \\
 0 & 1 & -1/2 & & 0& 1/2 & 0 \\ 
 0 & 5 & -2 & & -2 & 0 & 1 \\
\end{pmatrix}"

"R_3=R_3-5R_2"


"\begin{pmatrix}
 1 & 0 & 1/2 & & 1& 1/2 & 0 \\
 0 & 1 & -1/2 & & 0& 1/2 & 0 \\ 
 0 & 0 & 1/2 & & -2 & -5/2 & 1 \\
\end{pmatrix}"

"R_3=2R_3"


"\begin{pmatrix}
 1 & 0 & 1/2 & & 1& 1/2 & 0 \\
 0 & 1 & -1/2 & & 0& 1/2 & 0 \\ 
 0 & 0 & 1 & & -4 & -5 & 2 \\
\end{pmatrix}"

"R_1=R_1-R_3\/2"


"\begin{pmatrix}
 1 & 0 & 0 & & 3 & 3 & -1 \\
 0 & 1 & -1/2 & & 0& 1/2 & 0 \\ 
 0 & 0 & 1 & & -4 & -5 & 2 \\
\end{pmatrix}"

"R_2=R_2+R_3\/2"


"\begin{pmatrix}
 1 & 0 & 0 & & 3 & 3 & -1 \\
 0 & 1 & 0 & & -2 & -2 & 1 \\ 
 0 & 0 & 1 & & -4 & -5 & 2 \\
\end{pmatrix}"

We are done. On the left is the identity matrix. On the right is the inverse matrix.


"A^{-1}=\begin{pmatrix}
 3 & 3 & -1 \\
 -2 & -2 & 1 \\ 
 -4 & -5 & 2 \\
\end{pmatrix}"

"X=A^{-1}B"

"=\begin{pmatrix}
 3 & 3 & -1 \\
 -2 & -2 & 1 \\ 
 -4 & -5 & 2 \\
\end{pmatrix}\begin{pmatrix}
 1\\ 
1\\
1
\end{pmatrix}"

"=\begin{pmatrix}
 3+3-1\\ 
-2-2+1\\
-4-5+2
\end{pmatrix}=\begin{pmatrix}
 5\\ 
-3\\
-7
\end{pmatrix}"

"(5, -3, -7)"